Consider a locus with two alleles - A and a. These alleles are codominant, meaning that the fitness of the heterozygote is halfway between either homozygote. Consider further a population of randomly mating green frogs where the genotype counts are AA = 500, Aa = 250, and aa = 250. In this population the relative fitnesses of each genotype are AA = 1.00, Aa = 0.80, and aa = 0.60. What is the mean relative fitness within this population? Please give your answer to two decimal places.
Answer #1

Mean fitness = summation (genotypic frequency × fitness)

Genotypic frequency -

AA = 500/1000 = 0.5

Aa = 250/1000 = 0.25

aa = 250/1000 = 0.25

Mean fitness = (0.5 × 1) + (0.25 × 0.8) + (0.25 × 0.6) = 0.5 + 0.2 + 0.15

= 0.85

Mean fitness is 0.85.

Please rate high.